# On the Capacity of Self-Attention

### v1

@@ Appendix D Related Work
A growing body of work quantifies how many input–label associations Transformers—and, more narrowly, the attention mechanism—can memorize. As described in Section 2, the bounds from the memorization setting do not directly imply bounds on RGR, nor the other way around. For the attention module itself, [42] prove that a single MHA layer with $h$ heads can memorize $\Omega\!\big(h\,\min\{\ell,d_{k}\}\big)$ examples under a linear-independence assumption on the inputs, highlighting linear scaling in $h$ and the role of the per-head key/query width $d_{k}$. Complementary analyses bound attention’s memory depth and clarify depth–capacity trade-offs [41]. Moving to full Transformers, constructive results show that (under token-wise $(r,\delta)$-separated inputs) a stack of $2\ell$ self-attention layers suffices to memorize $N$ sequences with $\tilde{O}\!\big(\ell+\sqrt{\ell N}\big)$ parameters [35]; even a single-layer, single-head Transformer has nontrivial capacity under the same separatedness assumption, whereas replacing softmax by hardmax breaks memorization [31].
Beyond construction-style bounds, [41] give general upper and lower bounds for next-token prediction that scale as $\Theta(\omega N)$ in the presence of positional encodings and a vocabulary of size $\omega$, and [8] show that a single-layer Transformer can memorize when sequences are sufficiently zero-padded (though not in a parameter-optimal way). Classical results for ReLU networks connect parameter counts to memorization thresholds and VC-style capacity [58, 59]. More closely related to our focus on resource efficiency, [32] establish nearly matching upper/lower bounds on the minimal parameter count needed for memorization in Transformers: $\tilde{O}(\sqrt{N})$ parameters are sufficient (and necessary up to logs) for next-token prediction, and $\tilde{O}(\sqrt{\ell N})$ for sequence-to-sequence, under token-wise separatedness; they further suggest that self-attention effectively identifies sequences while the feed-forward network can become the bottleneck when associating labels.
@@
A concurrent line of work analyzes how networks compute many features in superposition, with lower and upper bounds for narrow MLPs and constructive designs for multi-feature computation [2, 1, 26]. The capacity limits shown there (stated as upper and lower bounds on neurons to compute a number of Boolean functions in parallel) are complementary to ours in terms of architectures: their focus is on MLPs while ours is on attention.
Foundational depth-separation and minimal-width universality results motivate the proof template we adopt—information-theoretic lower bounds matched by explicit constructions [55, 25, 34, 12]. In attention, constructive correspondences also explain how multi-head architectures partition pattern spaces; e.g., with relative positions, $s^{2}$ heads can realize any $s\times s$ convolution [10]. Our constructions similarly partition relational signal across heads to mitigate interference when $d_{\text{model}}\!\ll\!m$, explaining the empirical advantage of many small heads and clarifying when too-small $d_{k}$ triggers low-rank failure [5].
@@
A growing body of theory isolates how dimensional resources govern attention’s representational power. Universality guarantees establish that sufficiently resourced Transformers can approximate sequence-to-sequence functions [70], while more refined results show task-dependent strengths and weaknesses [52]. Focusing on the attention map, [39] prove that with fixed error and sparsity, self-attention can approximate dynamic sparse right-stochastic matrices using only $O(\log\ell)$ hidden dimensions (for context length $\ell$), echoing the role of near-orthogonality we exploit in our constructions. Conversely, [5] identify a per-head low-rank bottleneck: when $d_{k}<\ell$, a head cannot realize arbitrary $\ell\times\ell$ stochastic attention matrices. This clarifies a trade-off inside the total key/query budget $D_{K}=h\,d_{k}$: pushing $D_{K}$ into many tiny heads can induce head-wise rank limits.
Beyond these, several works develop structural and inductive-bias characterizations of self-attention. [15] show that pure attention without mixing loses rank doubly-exponentially with depth, explaining failure modes in deep attention stacks and underscoring the role of residual mixing. [17] analyze variable creation and sparsity patterns induced by softmax, while [51] use convex duality to give optimization- and geometry-based interpretations of ViT attention. For sample complexity and approximation, [37] study learning and generalization of shallow ViTs; rates and approximation guarantees have been developed for Transformer encoders and sequence models [23, 54, 30]. Recent generalization bounds that are (largely) sequence-length independent sharpen this picture [56]. Finally, theory has also pinpointed sparsity-oriented inductive biases: Transformers provably learn sparse token selection that FCNs cannot [68], and exhibit a simplicity bias for sparse Boolean functions [4].
Empirical observations likewise single out the key/query channel as an operative budget. Our results formalize this perspective for a concrete family (RGR), deriving matching lower and constructive upper bounds in terms of $D_{K}$ and the number of relations.
@@
Formal-language analyses delimit what fixed-size attention can recognize. Beyond general universality [70], there are sharp impossibility results for periodic and hierarchical languages [24, 3, 69]. Recent work uses communication-complexity arguments to show single-layer self-attention struggles with function composition at fixed embedding/heads, e.g., “grandparent-of” requires resources that scale with domain size [47]. Complementing these, [40] identify additional structural constraints on what Transformers can compute under realistic resource regimes. We view these results as orthogonal to RGR: they characterize classes of computations, whereas we fix a relational family and ask how much key/query budget is necessary and sufficient to represent its edges across arbitrary contexts.
In‑Context Learning and Algorithmic Views of Self‑Attention
A complementary line of theory frames Transformers—and attention in particular—as executing algorithms over the context. [38] analyze generalization and implicit model selection in in‑context learning; [65] give evidence that Transformers can implement gradient-descent‑like updates in context; and [20] characterize which simple function classes are learnable in context. These works clarify how attention can implement algorithmic behaviors over token sets, while our RGR focus quantifies the key–query capacity required to retrieve relational edges reliably.
@@
An alternative way to constrain attention is by controlling the connectivity pattern of the attention graph. Even $O(\ell)$-sparse patterns can be universal under appropriate designs [71], and systematic pruning of dense patterns maps out cost–performance frontiers [67]. Our analysis treats connectivity as not the bottleneck: given the ability to attend broadly, the limiting factor for RGR is how much relational information can be encoded and separated in keys/queries as $m$ and $m^{\prime}$ grow.
@@
Mechanistic interpretability consistently finds that specific heads specialize to linguistic relations [9, 63]. At the same time, many trained heads can be pruned with small accuracy loss [45, 64], indicating redundancy. Information-bottleneck analyses at the head/layer level quantify such redundancy and attribution in both language and vision models [49, 27], and architectural proposals target representation bottlenecks [21]. Our results supply a capacity-theoretic backbone for these observations: for RGR, performance transitions are governed primarily by $D_{K}$; distributing $D_{K}$ across heads reduces interference between superposed relations, but overly small $d_{k}$ per head incurs rank limits—predicting both specialization and safe pruning regimes.
@@
Modern Hopfield networks are equivalent, in a precise sense, to attention updates and can store exponentially many patterns in the associative dimension with single-step retrieval [50]. FFN layers in Transformers have also been interpreted as key–value memories [22]. Our results complement this memory-centric view by isolating the addressing budget: how much key/query capacity is required to select the correct neighbors (edges) for arbitrary contexts. Together these views separate storage capacity from the cost of accurate retrieval/selection in the key–query channel.
@@
Expressivity of graph Transformers is shaped by structural encodings and higher-order tokenization. SEG-WL analyses show that structural features (e.g., SPIS encodings) set the attainable expressivity ceiling and can be matched by simple Transformer variants [75]. Higher-order graph Transformers reach (or fall short of) $t$-WL power depending on whether explicit tuple indices and structural signals are provided [74]. In the vision setting, [29] prove that ViTs can learn spatial structure under appropriate conditions, resonating with our assumptions that near-orthogonal embeddings and structural signals determine how efficiently edges can be packed and recovered; given such signals, our $D_{K}$-based bounds become tight predictors of success.

### v3

@@ Appendix D Related Work
A growing body of work quantifies how many input–label associations Transformers—and, more narrowly, the attention mechanism—can memorize. As described in Section 2, the bounds from the memorization setting do not directly imply bounds on RGR, nor the other way around. For the attention module itself, Mahdavi et al. (2024) prove that a single MHA layer with $h$ heads can memorize $\Omega\!\big(h\,\min\{\ell,d_{k}\}\big)$ examples under a linear-independence assumption on the inputs, highlighting linear scaling in $h$ and the role of the per-head key/query width $d_{k}$. Complementary analyses bound attention’s memory depth and clarify depth–capacity trade-offs Madden et al. (2025). Moving to full Transformers, constructive results show that (under token-wise $(r,\delta)$-separated inputs) a stack of $2\ell$ self-attention layers suffices to memorize $N$ sequences with $\tilde{O}\!\big(\ell+\sqrt{\ell N}\big)$ parameters Kim et al. (2023); even a single-layer, single-head Transformer has nontrivial capacity under the same separatedness assumption, whereas replacing softmax by hardmax breaks memorization Kajitsuka and Sato (2024).
Beyond construction-style bounds, Madden et al. (2025) give general upper and lower bounds for next-token prediction that scale as $\Theta(\omega N)$ in the presence of positional encodings and a vocabulary of size $\omega$, and Chen and Zou (2024) show that a single-layer Transformer can memorize when sequences are sufficiently zero-padded (though not in a parameter-optimal way). Classical results for ReLU networks connect parameter counts to memorization thresholds and VC-style capacity Vardi et al. (2020, 2021). More closely related to our focus on resource efficiency, Kajitsuka and Sato (2025) establish nearly matching upper/lower bounds on the minimal parameter count needed for memorization in Transformers: $\tilde{O}(\sqrt{N})$ parameters are sufficient (and necessary up to logs) for next-token prediction, and $\tilde{O}(\sqrt{\ell N})$ for sequence-to-sequence, under token-wise separatedness; they further suggest that self-attention effectively identifies sequences while the feed-forward network can become the bottleneck when associating labels.
@@
A concurrent line of work analyzes how networks compute many features in superposition, with lower and upper bounds for narrow MLPs and constructive designs for multi-feature computation Adler and Shavit (2024); Adler et al. (2025); Hänni et al. (2024). Recent work further connects superposition to robust scaling behavior, providing additional evidence that feature packing can underwrite smooth scaling trends across model sizes Liu et al. (2025). The capacity limits shown in this line are complementary to ours in terms of architectures: their focus is on MLPs while ours is on attention and, in particular, on the key–query channel.
Foundational depth-separation and minimal-width universality results motivate the proof template we adopt—information-theoretic lower bounds matched by explicit constructions Telgarsky (2016); Hanin and Sellke (2019); Kidger and Lyons (2020); Cybenko (1989). In attention, constructive correspondences also explain how multi-head architectures partition pattern spaces; e.g., with relative positions, $s^{2}$ heads can realize any $s\times s$ convolution Cordonnier et al. (2020b). Our constructions similarly partition relational signal across heads to mitigate interference when $d_{\text{model}}\!\ll\!m$, explaining the empirical advantage of many small heads and clarifying when too-small $d_{k}$ triggers low-rank failure Bhojanapalli et al. (2020).
@@
A growing body of theory isolates how dimensional resources govern attention’s representational power. Universality guarantees establish that sufficiently resourced Transformers can approximate sequence-to-sequence functions Yun et al. (2020a), while more refined results show task-dependent strengths and weaknesses Sanford et al. (2023). Focusing on the attention map, Likhosherstov et al. (2021) prove that with fixed error and sparsity, self-attention can approximate dynamic sparse right-stochastic matrices using only $O(\log\ell)$ hidden dimensions (for context length $\ell$), echoing the role of near-orthogonality we exploit in our constructions. Conversely, Bhojanapalli et al. (2020) identify a per-head low-rank bottleneck: when $d_{k}<\ell$, a head cannot realize arbitrary $\ell\times\ell$ stochastic attention matrices. This clarifies a trade-off inside the total key/query budget $D_{K}=h\,d_{k}$: pushing $D_{K}$ into many tiny heads can induce head-wise rank limits.
Beyond these, several works develop structural and inductive-bias characterizations of self-attention. Dong et al. (2021) show that pure attention without mixing loses rank doubly-exponentially with depth, explaining failure modes in deep attention stacks and underscoring the role of residual mixing. Edelman et al. (2022) analyze variable creation and sparsity patterns induced by softmax, while Sahiner et al. (2022) use convex duality to give optimization- and geometry-based interpretations of ViT attention. For sample complexity and approximation, Li et al. (2023a) study learning and generalization of shallow ViTs; rates and approximation guarantees have been developed for Transformer encoders and sequence models Gurevych et al. (2022); Takakura and Suzuki (2023); Jiang and Li (2024). Recent generalization bounds that are (largely) sequence-length independent sharpen this picture Trauger and Tewari (2024). Finally, theory has also pinpointed sparsity-oriented inductive biases: Transformers provably learn sparse token selection that FCNs cannot Wang et al. (2024), and exhibit a simplicity bias for sparse Boolean functions Bhattamishra et al. (2023). Recent results show that when embeddings are learned, attention can provably focus on informative tokens under suitable conditions Wu et al. (2025).
Empirical observations likewise single out the key/query channel as an operative budget. Our results formalize this perspective for a concrete relational family (RGR), deriving matching lower and constructive upper bounds in terms of $D_{K}$ and the number of relations.
@@
Formal-language analyses delimit what fixed-size attention can recognize. Beyond general universality Yun et al. (2020a), there are sharp impossibility results for periodic and hierarchical languages Hahn (2020); Bhattamishra et al. (2020); Yang et al. (2024). Recent work uses communication-complexity arguments to show single-layer self-attention struggles with function composition at fixed embedding/heads, e.g., “grandparent-of” requires resources that scale with domain size Peng et al. (2024). Complementing these, Luo et al. (2022) identify additional structural constraints on what Transformers can compute under realistic resource regimes. We view these results as orthogonal to RGR: they characterize classes of computations, whereas we fix a relational family and ask how much key/query budget is necessary and sufficient to represent its edges across arbitrary contexts.
In-Context Learning and Algorithmic Views of Self-Attention
A complementary line of theory frames Transformers—and attention in particular—as executing algorithms over the context. Li et al. (2023b) analyze generalization and implicit model selection in in-context learning; von Oswald et al. (2022) give evidence that Transformers can implement gradient-descent-like updates in context; and Garg et al. (2022) characterize which simple function classes are learnable in context. Recent theory also isolates the role of positional information in algorithmic execution: Back de Luca et al. (2025) study positional attention, where attention weights depend only on positional encodings, and show such models can retain strong expressivity for parallel algorithmic computation (with depth/sample-complexity tradeoffs). These works clarify how attention can implement algorithmic behaviors, while our RGR focus quantifies the key–query capacity required to retrieve relational edges reliably.
@@
An alternative way to constrain attention is by controlling the connectivity pattern of the attention graph. Even $O(\ell)$-sparse patterns can be universal under appropriate designs Yun et al. (2020b), and systematic pruning of dense patterns maps out cost–performance frontiers Wang et al. (2022). Our analysis treats connectivity as not the bottleneck: given the ability to attend broadly, the limiting factor for RGR is how much relational information can be encoded and separated in keys/queries as $m$ and $m^{\prime}$ grow.
Emergent Sparsity and Transition Phenomena in Attention
Beyond hard-wiring sparsity, several recent works study how sparse (or structured) attention patterns emerge during training and how these transitions depend on data statistics. For example, Zucchet et al. (2025) analyze when sparse attention arises and how repetition and distributional structure can accelerate its emergence. This line is complementary to our setting: we characterize budget-driven transitions in relational recovery as $D_{K}$ varies under controlled relational structure (RGR), whereas emergence work focuses on training dynamics and when attention becomes sparse or selective.
@@
Mechanistic interpretability consistently finds that specific heads specialize to linguistic relations Clark et al. (2019); Vig and Belinkov (2019). At the same time, many trained heads can be pruned with small accuracy loss Michel et al. (2019); Voita et al. (2019), indicating redundancy. Information-bottleneck analyses at the head/layer level quantify such redundancy and attribution in both language and vision models Qian et al. (2025); Hong et al. (2025), and architectural proposals target representation bottlenecks Gerasimov et al. (2025). More recently, causal attribution frameworks explicitly learn gates over heads to characterize facilitating vs. interfering roles and head interactions, reinforcing that head effects can be context-dependent and non-modular Nam et al. (2025). Our results supply a capacity-theoretic backbone for these observations: for RGR, performance transitions are governed primarily by $D_{K}$; distributing $D_{K}$ across heads reduces interference between superposed relations, but overly small $d_{k}$ per head incurs rank limits—predicting both specialization and safe pruning regimes.
@@
Modern Hopfield networks are equivalent, in a precise sense, to attention updates and can store exponentially many patterns in the associative dimension with single-step retrieval Ramsauer et al. (2021). FFN layers in Transformers have also been interpreted as key–value memories Geva et al. (2021). Our results complement this memory-centric view by isolating the addressing budget: how much key/query capacity is required to select the correct neighbors (edges) for arbitrary contexts. Together these views separate storage capacity from the cost of accurate retrieval/selection in the key–query channel.
A recent synthesis by Zhong et al. (2025) frames both self-attention and FFNs as instances of kernelized associative memory, and proposes a retrieval signal-to-noise ratio (SNR) to quantify recall fidelity as the number of stored key–value pairs grows. Their analysis highlights how the exponential kernel underlying softmax attention can dramatically improve retrieval SNR relative to linear (and ReLU-like) kernels, and connects kernel choice to a precision–superposition tradeoff that influences polysemanticity. This viewpoint is complementary to ours: whereas their capacity proxy is recall accuracy for stored associations under distributional assumptions, our RGR formulation studies worst-case relational addressing across all contexts and all graphs in a family, yielding necessary and sufficient bounds in terms of the total key dimension $D_{K}$.
@@
Expressivity of graph Transformers is shaped by structural encodings and higher-order tokenization. SEG-WL analyses show that structural features (e.g., SPIS encodings) set the attainable expressivity ceiling and can be matched by simple Transformer variants Zhu et al. (2023). Recent work has also aimed to bridge the theory–practice gap in graph Transformers by unifying architectural choices around attention and positional/structural encodings and validating them at scale Stoll et al. (2025). Higher-order graph Transformers reach (or fall short of) $t$-WL power depending on whether explicit tuple indices and structural signals are provided Zhou et al. (2024). In the vision setting, Jelassi et al. (2022) prove that ViTs can learn spatial structure under appropriate conditions, resonating with our assumptions that near-orthogonal embeddings and structural signals determine how efficiently edges can be packed and recovered; given such signals, our $D_{K}$-based bounds become tight predictors of success.
Linear / softmax-free attention.
A large line of work replaces or approximates softmax attention to reduce the quadratic cost in sequence length, often by expressing attention as a kernel feature-map product that can be evaluated associatively (so-called linear attention) Katharopoulos et al. (2020). Follow-up methods improve the fidelity of softmax approximations via random-feature estimators Choromanski et al. (2021); Peng et al. (2021) or exploit low-rank structure of the attention map Wang et al. (2020), while other work proposes alternative normalizers / reweightings that aim to recover some of softmax’s concentration behavior with linear-time computation Qin et al. (2022). Recent work continues to probe the softmax/linear divide from several angles: (i) gating-based operators reintroduce nonlinearity and sparsity while mitigating attention pathologies (e.g., attention sinks), highlighting the functional role of competitive nonlinearities in routing Qiu et al. (2025); (ii) feature-efficiency analyses ask how many linear-attention features are needed to distill or approximate softmax well under compute constraints Nishikawa et al. (2025); (iii) simplified statistical models and mean-field analyses provide theoretical accounts of when softmax enjoys intrinsic advantages and how it behaves during learning Duranthon et al. (2025); Dohmatob (2025); and (iv) asymptotic results identify regimes where softmax attention approaches linear behavior, clarifying when the distinction can blur Boursier and Boyer (2025). From a unifying associative-memory perspective, Zhong et al. (2025) analyze softmax vs. linear attention via retrieval-SNR and propose update-rule variants (e.g., delta-rule hybrids) that reintroduce selective overwrite/forgetting while retaining high-fidelity kernelized retrieval. Our focus is complementary: rather than proposing a new efficient operator, we use RGR to isolate how the presence and type of nonlinearity in the QK routing computation affects relational capacity and multi-head interference reduction, clarifying when purely linear score aggregation can lose the multi-head benefit and why even simple nonlinearities (e.g., max aggregation) can change this behavior.